JUST THE QUESTION 16 PLEASE THE FINAL PART OF C IS DEADWEIGHT LOSS AND COMPARE THEM WITH YOUR FINDINGS ON A

co(b) each firm produces (c) each firm is a price taker. (d) there are few firms in the market. le) each firm observes a horizo
Answer #1

9 -60-8 IP=60-0) [MR=60-20) IMC=AC=10) Monopoly profit is mak when MR-me Gorzo = lo @ = cle=35] Monopoly = 351257_10125)=625